# Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models

### v2

@@ 1 Introduction
Large Language Models (LLMs) have began a groundbreaking era in artificial general intelligence (AGI), demonstrating extraordinary capabilities across a wide range of domains that require intricate reasoning and specialized knowledge. These models excel in areas such as mathematical reasoning/problem solving (Cobbe et al., 2021; Wei et al., 2022; Lewkowycz et al., 2022), code generation/programming (Chen et al., 2021; Austin et al., 2021; Li et al., 2022), text generation (Bubeck et al., 2023; Anil et al., 2023; Touvron et al., 2023), summarization and creative writing, among others. A significant advancement in LLMs is the post-pretraining alignment with the more desirable behaviors (Mishra et al., 2021; Victor et al., 2022; Chung et al., 2022; Thoppilan et al., 2022), a process often reliant on the costly human-annotated data. Typical alignment methods include Supervised Fine-Tuning (SFT) (Ouyang et al., 2022; Tunstall et al., 2023a) based on human demonstrations, and Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Bai et al., 2022a) based on human preferences.
All the aforementioned alignment methods require a substantial volume of human annotated data. Therefore, there is increasing interest in developing fine-tuning methods that can effectively utilize human data, thereby streamlining the alignment process. This motivates us to study fine-tuning LLMs without the need for additional human-annotated data beyond the fine-tuning dataset. Our study is also related to the broader goal of converting weak models to strong models without the requirement for extra training data, which is of central interest in machine learning that can be traced back to the boosting algorithms (Kearns and Valiant, 1994; Schapire, 1990; Freund, 1995; Freund and Schapire, 1997). The self-training algorithm (Vapnik, 1999; Grandvalet and Bengio, 2004; Lee, 2013) has also been proved to be able to convert weak learners to strong learners in mixture models without the need for additional labeled data (Frei et al., 2022; Kou et al., 2022). However, the pursuit of autonomously enhancing a weak LLM without external guidance is both intriguing and understudied. This raises the following question:
@@
Notation.
We use lowercase letters and lowercase boldface letters to denote scalars and vectors, respectively. We use $[N]$ to denote the index set $\{1,\dots,N\}$. In the function space, let $\mathcal{F}$ be the function class. The symbol $q_{\mathrm{data}}$ designates the target data distribution, while $p$ represents the conditional probability of LLM’s response (i.e., LLM policy).
@@ 2 Related Work
Self-Play.
Self-play (Samuel, 1959; Tesauro et al., 1995), where the algorithm learns by playing against itself, has gained notable attention due to its effectiveness in multi-agent reinforcement learning (MARL). This method involves agents engaging in interactions with copies of themselves, enabling an increasing level of challenge and complexity within the learning environment. A fundamental work in the field of self-play is AlphaGo Zero (Silver et al., 2017b), which demonstrated exceptional performance against human players using a self-play learning scheme. Subsequent research has expanded upon the concept of self-play, exploring various adaptations and implementations (Anthony et al., 2017; Lanctot et al., 2017; Bansal et al., 2018; Hernandez-Leal et al., 2018; Muller et al., 2019; Vinyals et al., 2019). Our method takes the self-play approach akin to AlphaGo Zero, which can convert a weak model to a strong one without additional human-annotated data. While the effectiveness of self-play in MARL is well-established, to our knowledge, our work is the first to apply this approach to the enhancement of LLMs.
Synthetic Data for LLMs.
In the context of supervised fine-tuning (SFT) of LLMs, human-crafted data has proven to be a remarkably effective source that enhances the performance of LLMs on tasks such as code generation (Roziere et al., 2023; Yang et al., 2023) and mathematical reasoning (Yuan et al., 2023; Luo et al., 2023). While human data typically exhibits high quality, acquiring sufficient amount of such data poses a challenge in cost. In light of this consideration, the use of synthetic data has become increasingly popular and considered as a proxy for human data. This approach primarily leverages advanced LLMs such as the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) as the guidance to generate high-quality data (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023; Li et al., 2023). Recent research has also highlighted the rephrasing capability of LLMs in prompting for better LLM response (Deng et al., 2023; Prasad et al., 2023) as well as augmenting synthetic data for more effective SFT (Yu et al., 2023; Liu et al., 2023). In contrast to prior studies that utilized more advanced models for synthetic data generation when pre-training or fine-tuning a target model, our approach directly generates synthetic data from the target model itself.
@@ 3.2 RL Fine-Tuning
$L_{\mathrm{RL}}(\bm{\theta})=\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\bm{\theta}}(\cdot|\mathbf{x})}[r(\mathbf{x},\mathbf{y})]-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p_{\bm{\theta}}(\cdot|\mathbf{x})||p_{\mathrm{ref}}(\cdot|\mathbf{x})\big{)},$
@@ 4.1 Self-Play Fine-Tuning (SPIN)
We begin with illustrating how we expect a main player is trained to distinguish LLM responses from human responses. Motivated by integral probability metric (IPM) (Müller, 1997), we formulate our objective function such that the main player $f_{t+1}$ maximizes the expected value gap between the target data distribution $p_{\mathrm{data}}$ and the opponent player’s distribution $p_{\bm{\theta}_{t}}$:
$f_{t+1}$ $=\mathop{\mathrm{argmax}}_{f\in\mathcal{F}_{t}}\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\big{[}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{]},$ (4.1)
@@
where $\mathcal{F}_{t}$ is a sequence of highly expressive function classes that we will determine in later deduction. The subscript $t$ in $\mathcal{F}_{t}$ is due to that the function class is dependent on $p_{\bm{\theta}_{t}}$. Given such a $f_{t+1}$ and a response sequence $\mathbf{y}$ to the prompt $\mathbf{x}$, the value of $f_{t+1}(\mathbf{x},\mathbf{y})$ reflects the main player’s degree of belief that $\mathbf{y}$ originates from $p_{\mathrm{data}}$ rather than $p_{\bm{\theta}_{t}}$. Ideally, the main player $f_{t+1}$ should yield a high value when $\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x})$ and a low value when $\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, where $p_{\bm{\theta}_{t}}$ is the opponent’s distribution. Instead of solving (4.1), we can also solve the following more general optimization problem,
$f_{t+1}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}_{t}}\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),y^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\big{[}\ell\big{(}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{)}\big{]},$ (4.2)
@@
where $\ell(\cdot)$ is a loss function that is both monotonically decreasing and convex. For example, a linear loss function $\ell(t)=-t$ reduces (4.2) to the minimization version of (4.1). However, the use of a linear loss function results in an unbounded objective value, which, during continuous training, leads to a negative infinite value of $f(\mathbf{x},\mathbf{y}^{\prime})$ on the opponent player’s responses. Therefore, in our work, we choose the logistic loss function $\ell(t):=\log(1+\exp(-t))$ for its non-negativity, smoothness, and exponentially decaying tail as $t\rightarrow\infty$. Such a choice of loss function aids in preventing the excessive growth in the absolute value of $f$.
Updating the Opponent Player.
@@
Previously we have discussed the training of $f_{t+1}$ given the opponent player’s distribution $p_{\bm{\theta}_{t}}$. Now suppose we have optimized our main player $f_{t+1}$ that can distinguish $p_{\mathrm{data}}$ from $p_{\bm{\theta}_{t}}$, within a certain function class $\mathcal{F}_{t}$, we elaborate how we get parameter $\bm{\theta}_{t+1}$ of the opponent player. Specifically, when presented with two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ to the same prompt $\mathbf{x}$, $f_{t+1}$ assesses the values $f_{t+1}(\mathbf{x},\mathbf{y})$ and $f_{t+1}(\mathbf{x},\mathbf{y}^{\prime})$. It then infers that the response with the higher value is from the real data distribution $p_{\mathrm{data}}$ and the response with lower value is attributed to the LLM $p_{\bm{\theta}_{t}}$. Subsequently, the objective of the opponent player is to find a better LLM that generates responses indistinguishable from $p_{\mathrm{data}}$ for the main player. This is achieved by maximizing the expected value $\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p(\cdot|\mathbf{x})}[f_{t+1}(\mathbf{x},\mathbf{y})]$. In addition, to prevent excessive deviation of $p_{\bm{\theta}_{t+1}}$ from $p_{\bm{\theta}_{t}}$ and stabilize the self-play, we incorporate a Kullback-Leibler (KL) regularization term. Putting these together gives rise to the following optimization problem:
$\mathop{\mathrm{argmax}}_{p}\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p(\cdot|\mathbf{x})}[f_{t+1}(\mathbf{x},\mathbf{y})]-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p(\cdot|\mathbf{x})||p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})\big{)},$ (4.3)
@@
End-to-end Training Objective.
We integrate the previously discussed two steps into a single end-to-end training objective with an update rule of $\bm{\theta}_{t+1}$. Specifically, plugging (4.5) into (4.2) arrives at the update rule $\bm{\theta}_{t+1}=\mathop{\mathrm{argmin}}_{\bm{\theta}\in\bm{\Theta}}L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t})$, where $L_{\texttt{SPIN}}$ is the training objective defined as follows
$L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t})=\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\bigg{[}\ell\bigg{(}\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}-\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}^{\prime}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x})}\bigg{)}\bigg{]}.$ (4.7)
We summarize the iterative self-play process of our method SPIN as follows,
$\ldots\quad\rightarrow\underbrace{p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}_{\text{Opponent Player at $t$}}\rightarrow\quad\underbrace{\lambda\cdot\log\frac{p_{\bm{\theta}_{t+1}}(\cdot|\mathbf{x})}{p_{\mathrm{\bm{\theta}_{t}}}(\cdot|\mathbf{x})}}_{\text{Main Player at $t+1$}}\quad\rightarrow\underbrace{p_{\bm{\theta}_{t+1}}(\cdot|\mathbf{x})}_{\text{Opponent Player at $t+1$}}\rightarrow\quad\ldots$
@@
Remark 4.1.
(4.7) bears resemblance to direct preference optimization (DPO) (Rafailov et al., 2023) for RL fine-tuning. However, SPIN exhibits significant distinctions with DPO. Specifically, SPIN is applied to supervised fine-tuning (SFT) and relies solely on the SFT dataset, represented by pairs $(\mathbf{x},\mathbf{y})$. In sharp contrast, DPO is designed for RL fine-tuning and necessitates a preference dataset, represented by $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote the winner (chosen) and loser (rejected) responses, respectively. DPO demands that, at the instance level, $\mathbf{y}_{w}$ is superior to $\mathbf{y}_{l}$. In comparison, our method requires that, at the distribution level, the target $p_{\mathrm{data}}$ should be distinguishable from the weak LLM $p_{\bm{\theta}}$ before it becomes a strong one. In terms of algorithm design, DPO implements a single-iteration approach, while our method facilitates an iterative self-play strategy, as outlined in Algorithm 1.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have began a groundbreaking era in artificial general intelligence (AGI), demonstrating extraordinary capabilities across a wide range of domains that require intricate reasoning and specialized knowledge. These models excel in areas such as mathematical reasoning/problem solving (Cobbe et al., 2021; Wei et al., 2022; Lewkowycz et al., 2022), code generation/programming (Chen et al., 2021; Austin et al., 2021; Li et al., 2022), text generation (Bubeck et al., 2023; Anil et al., 2023; Touvron et al., 2023), summarization and creative writing, among others. A significant advancement in LLMs is the post-pre-training alignment with the more desirable behaviors (Mishra et al., 2021; Victor et al., 2022; Chung et al., 2022; Thoppilan et al., 2022), a process often reliant on the costly human-annotated data. Typical alignment methods include Supervised Fine-Tuning (SFT) (Ouyang et al., 2022; Tunstall et al., 2023a) based on human demonstrations, and Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Bai et al., 2022a) based on human preferences.
All the aforementioned alignment methods require a substantial volume of human annotated data. Therefore, there is increasing interest in developing fine-tuning methods that can effectively utilize human data, thereby streamlining the alignment process. This motivates us to study fine-tuning LLMs without the need for additional human-annotated data beyond the fine-tuning dataset. Our study is also related to the broader goal of converting weak models to strong models without the requirement for extra training data, which is of central interest in machine learning that can be traced back to the boosting algorithms (Kearns & Valiant, 1994; Schapire, 1990; Freund, 1995; Freund & Schapire, 1997). The self-training algorithm (Vapnik, 1999; Grandvalet & Bengio, 2004; Lee, 2013) has also been proved to be able to convert weak learners to strong learners in mixture models without the need for additional labeled data (Frei et al., 2022; Kou et al., 2022). However, the pursuit of autonomously enhancing a weak LLM without external guidance is both intriguing and understudied. This raises the following question:
@@
Notation. We use lowercase letters and lowercase boldface letters to denote scalars and vectors, respectively. We use $[N]$ to denote the index set $\{1,\dots,N\}$. In the function space, let $\mathcal{F}$ be the function class. The symbol $q_{\mathrm{data}}$ designates the target data distribution, while $p$ represents the conditional probability of LLM’s response (i.e., LLM policy).
@@ 2 Related Work
Self-Play. Self-play (Samuel, 1959; Tesauro et al., 1995), where the algorithm learns by playing against itself, has gained notable attention due to its effectiveness in multi-agent reinforcement learning (MARL). This method involves agents engaging in interactions with copies of themselves, enabling an increasing level of challenge and complexity within the learning environment. A fundamental work in the field of self-play is AlphaGo Zero (Silver et al., 2017b), which demonstrated exceptional performance against human players using a self-play learning scheme. Subsequent research has expanded upon the concept of self-play, exploring various adaptations and implementations (Anthony et al., 2017; Lanctot et al., 2017; Bansal et al., 2018; Hernandez-Leal et al., 2018; Muller et al., 2019; Vinyals et al., 2019). Our method takes the self-play approach akin to AlphaGo Zero, which can convert a weak model to a strong one without additional human-annotated data. While the effectiveness of self-play in MARL is well-established, to our knowledge, our work is the first to apply this approach to the enhancement of LLMs.
Synthetic Data for LLMs. In the context of supervised fine-tuning (SFT) of LLMs, human-crafted data has proven to be a remarkably effective source that enhances the performance of LLMs on tasks such as code generation (Roziere et al., 2023; Yang et al., 2023) and mathematical reasoning (Yuan et al., 2023; Luo et al., 2023). While human data typically exhibits high quality, acquiring sufficient amount of such data poses a challenge in cost. In light of this consideration, the use of synthetic data has become increasingly popular and considered as a proxy for human data. This approach primarily leverages advanced LLMs such as the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) as the guidance to generate high-quality data (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023; Li et al., 2023). Recent research has also highlighted the rephrasing capability of LLMs in prompting for better LLM response (Deng et al., 2023; Prasad et al., 2023) as well as augmenting synthetic data for more effective SFT (Yu et al., 2023; Liu et al., 2023). In contrast to prior studies that utilized more advanced models for synthetic data generation when pre-training or fine-tuning a target model, our approach directly generates synthetic data from the target model itself.
@@ 3.2 RL Fine-Tuning
$L_{\mathrm{RL}}(\bm{\theta})$ $=\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\bm{\theta}}(\cdot|\mathbf{x})}[r(\mathbf{x},\mathbf{y})]$
$\qquad-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p_{\bm{\theta}}(\cdot|\mathbf{x})||p_{\mathrm{ref}}(\cdot|\mathbf{x})\big{)},$
@@ 4.1 Self-Play Fine-Tuning (SPIN)
Training the Main Player. We begin with illustrating how we expect a main player is trained to distinguish LLM responses from human responses. Motivated by integral probability metric (IPM) (Müller, 1997), we formulate our objective function such that the main player $f_{t+1}$ maximizes the expected value gap between the target data distribution $p_{\mathrm{data}}$ and the opponent player’s distribution $p_{\bm{\theta}_{t}}$:
$f_{t+1}$ $=\mathop{\mathrm{argmax}}_{f\in\mathcal{F}_{t}}\mathbb{E}\big{[}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{]},$ (4.1)
@@
where the expectation is computed over the distributions $\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, and $\mathcal{F}_{t}$ is a sequence of highly expressive function classes that we will determine in later deduction. The subscript $t$ in $\mathcal{F}_{t}$ is due to that the function class is dependent on $p_{\bm{\theta}_{t}}$. Given such a $f_{t+1}$ and a response sequence $\mathbf{y}$ to the prompt $\mathbf{x}$, the value of $f_{t+1}(\mathbf{x},\mathbf{y})$ reflects the main player’s degree of belief that $\mathbf{y}$ originates from $p_{\mathrm{data}}$ rather than $p_{\bm{\theta}_{t}}$. Ideally, the main player $f_{t+1}$ should yield a high value when $\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x})$ and a low value when $\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, where $p_{\bm{\theta}_{t}}$ is the opponent’s distribution. Instead of solving (4.1), we can also solve the following more general optimization problem,
$f_{t+1}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}_{t}}\mathbb{E}\big{[}\ell\big{(}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{)}\big{]},$ (4.2)
@@
where the expectation is computed over the distribution $\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),y^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$, and $\ell(\cdot)$ is a loss function that is both monotonically decreasing and convex. For example, a linear loss function $\ell(t)=-t$ reduces (4.2) to the minimization version of (4.1). However, the use of a linear loss function results in an unbounded objective value, which, during continuous training, leads to a negative infinite value of $f(\mathbf{x},\mathbf{y}^{\prime})$ on the opponent player’s responses. Therefore, in our work, we choose the logistic loss function $\ell(t):=\log(1+\exp(-t))$ for its non-negativity, smoothness, and exponentially decaying tail as $t\rightarrow\infty$. Such a choice of loss function aids in preventing the excessive growth in the absolute value of $f$.
@@
Updating the Opponent Player. Previously we have discussed the training of $f_{t+1}$ given the opponent player’s distribution $p_{\bm{\theta}_{t}}$. Now suppose we have optimized our main player $f_{t+1}$ that can distinguish $p_{\mathrm{data}}$ from $p_{\bm{\theta}_{t}}$, within a certain function class $\mathcal{F}_{t}$, we elaborate how we get parameter $\bm{\theta}_{t+1}$ of the opponent player. Specifically, when presented with two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ to the same prompt $\mathbf{x}$, $f_{t+1}$ assesses the values $f_{t+1}(\mathbf{x},\mathbf{y})$ and $f_{t+1}(\mathbf{x},\mathbf{y}^{\prime})$. It then infers that the response with the higher value is from the real data distribution $p_{\mathrm{data}}$ and the response with lower value is attributed to the LLM $p_{\bm{\theta}_{t}}$. Subsequently, the objective of the opponent player is to find a better LLM that generates responses indistinguishable from $p_{\mathrm{data}}$ for the main player. This is achieved by maximizing the expected value $\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p(\cdot|\mathbf{x})}[f_{t+1}(\mathbf{x},\mathbf{y})]$. In addition, to prevent excessive deviation of $p_{\bm{\theta}_{t+1}}$ from $p_{\bm{\theta}_{t}}$ and stabilize the self-play, we incorporate a Kullback-Leibler (KL) regularization term. Putting these together gives rise to the following optimization problem:
$\mathop{\mathrm{argmax}}_{p}\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p(\cdot|\mathbf{x})}[f_{t+1}(\mathbf{x},\mathbf{y})]$
$\qquad-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p(\cdot|\mathbf{x})||p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})\big{)},$ (4.3)
@@
End-to-end Training Objective. We integrate the previously discussed two steps into a single end-to-end training objective with an update rule of $\bm{\theta}_{t+1}$. Specifically, plugging (4.5) into (4.2) arrives at the update rule $\bm{\theta}_{t+1}=\mathop{\mathrm{argmin}}_{\bm{\theta}\in\bm{\Theta}}L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t})$, where $L_{\texttt{SPIN}}$ is the training objective defined as follows
$L_{\texttt{SPIN}}=\mathbb{E}\bigg{[}\ell\bigg{(}\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}-\lambda\log\frac{p_{\bm{\theta}}(\mathbf{y}^{\prime}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x})}\bigg{)}\bigg{]},$ (4.7)
where the expectation is computed over the distribution $\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),\mathbf{y}^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})$. We summarize the iterative self-play process of our method SPIN as follows,
$\ldots$ $\quad\rightarrow\underbrace{p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}_{\text{Opponent Player at $t$}}\rightarrow\quad\underbrace{\lambda\cdot\log\frac{p_{\bm{\theta}_{t+1}}(\cdot|\mathbf{x})}{p_{\mathrm{\bm{\theta}_{t}}}(\cdot|\mathbf{x})}}_{\text{Main Player at $t+1$}}$
$\quad\rightarrow\underbrace{p_{\bm{\theta}_{t+1}}(\cdot|\mathbf{x})}_{\text{Opponent Player at $t+1$}}\rightarrow\quad\ldots$
@@
4.2 Comparison between SPIN and DPO
In Section 4.1, we propose Self-Play Fine-Tuning (SPIN) with an end-to-end training objective (4.7) for each iteration. (4.7) bears resemblance to direct preference optimization (DPO) (Rafailov et al., 2023) for RL fine-tuning. However, SPIN and DPO are fundamentally different. DPO is based on the Bradley-Terry (BT) model: $p(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})=\frac{\exp(r^{*}(\mathbf{x},\mathbf{y}_{1}))}{\exp(r^{*}(\mathbf{x},\mathbf{y}_{1}))+\exp(r^{*}(\mathbf{x},\mathbf{y}_{2}))}$, and maximizes the log-likelihood of $p(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})$ by direct policy optimization without explicit reward estimation. In contrast, SPIN relies on maximizing the IPM to compete with an increasingly stronger version of itself. More detailed comparisons are highlighted as follows:
  1. 1.
    DPO does not inherently lead to iterative training. More specifically, DPO aims to match the preference probability $p(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})$ induced from its reward model with the data distribution $p_{\mathrm{data}}(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})$ in a single iteration. On the contrary, SPIN’s self-play mechanism naturally leads to an iterative training procedure. SPIN iteratively refines its generation distribution $p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})$ to match the target distribution $p_{\text{data}}(\mathbf{y}|\mathbf{x})$ across iterations.
  2. 2.
    SPIN only requires the SFT dataset, represented by pairs $(\mathbf{x},\mathbf{y})$. In contrast, DPO necessitates a preference dataset, represented by $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$, where $\mathbf{y}_{w}$ and $\mathbf{y}_{l}$ denote the winner (chosen) and loser (rejected) responses, respectively. Moreover, SPIN can be applied between SFT and RL fine-tuning.
  3. 3.
    In SPIN, we can choose different loss functions $\ell$ which only need to be convex and decreasing (detailed later in Theorem 5.2), which includes correlation loss, hinge loss, and logistic loss. Only when $\ell$ is chosen as the logistic loss would the training objective of SPIN become similar to that of DPO.
Recently, Xu et al. (2023) proposed to use iterative preference optimization with the Pairwise Cringe Loss (PCO), and generalized DPO to iterative DPO. Concurrent to our work, Yuan et al. (2024) further proposed a framework named “self-rewarding language models”, which leverages the LLM itself as the reward model to provide the preference feedback, and employs iterative DPO to train the LLM. Compared with Xu et al. (2023); Yuan et al. (2024), SPIN’s self-assessment is implicit, as no intermediate reward or preference feedback is required.